OpenStudy (anonymous):

A 208g oscillator has a speed of 112cm/s when its displacement is 1.60cm and 48.5cm/s when its displacement is 5.70cm .What is the oscillator's maximum speed?

OpenStudy (alekos):

is there more info to this question? i don't think there's enough to go on

OpenStudy (anonymous):

For a linear osscilator F = k*x dF/dt = k (dx/dt) Force = 0 at x=0 = maximum at x=maximum Force varies from zero to a maximum in a linear fashion so the velocity varies from zero to a maximum in a linear fashion. We need to find the parameters for velocity = slope*(displacement) + intercept y = mx + b Change between the two given points is velocity = (48.5-112) = -63.5 cm/s distance = (5.7-1.6) = 4.1 cm so slope = -(63.5/4.1) = -15.48 (cm/s)/(cm) y = (-15.48)*x + b using 112 = (-15.48)*1.6 + b gives b = 112 + (15.48*1.6) = 136.8 Maximum speed occurs at x=0 (the y-intercept) y = (-15.48)*0 + 136.8 = 136.8 cm/s Maximum displacement occurs at y = 0 = (-15.48)*x + 136.8 x = (136.8/15.48) = 8.83 cm

OpenStudy (alekos):

the formula your using is incorrect For displacement x = \[Asin \omega\] t where A is the max displacement \[\omega \] is the frequency and t is the time For velocity dx/dt = \[A \omega \cos \omega t\]